How much work must be done to accelerate an 800kg car from 15m/s to 30m/s?


1
Expert's answer
2022-01-10T16:56:15-0500

The energy-work theorem gives

"\Delta E=W\\
\frac{mv_2^2}{2}-\frac{mv_1^2}{2}=W""W=\rm\frac{800\:kg*(30\:m/s)^2}{2}-\rm\frac{800\:kg*(15\:m/s)^2}{2}"

"W=\rm 270\: 000\; J"